Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
14x + 18y = -54
-4x – 14y = 42

Answer by stanbon(75887)

14x + 18y = -54
-4x – 14y = 42
-----
Multiply thru 1st by 2:
Multiply thru 2nd by 7:
---
28x + 36y = -108
-28x -98y = 294
-----
Add and solve for "y":
62y = 186
y = 3
-------
Solve for "x":
-4x - 14y = 42
-4x - 14*3 = 42
----
-4x - 42 = 42
-4x = 84
x = -21
===================
Cheers,
Stan H.
========================